I am having trouble with this question:
Write a formula for the general term (the nth term) of the arithmetic sequence whose second term, a base7 is 4 and whose ninth term, a base is 12.
I am thinking the formula to use is an= a1+(n-1)d
an=1(12-1)4
Is this right?
If you could help me I would appreciate it


Answer by stanbon(75887)

Write a formula for the general term (the nth term) of the arithmetic sequence whose second term, a base7 is 4 and whose ninth term, a base is 12.
----
2nd term: a + (2-1)d = 4
9th term: a + (9-1)d = 12
------
Modify:
a + d = 4
a + 8d=12
====
Subtract and solve for "d":
7d = 8
d = 8/7
----
Solve for "a":
a+d = 4
a + (8/7) = 28/7
a = 20/7
=======
General term:
a(n) = (20/7)+(n-1)(8/7)
=========================
Cheers,
Stan H.